# Two-Person Interaction Augmentation with Skeleton Priors

### 08 Apr 2024 v1

@@ 3 Methodology
Figure 1: Overview of our model. The key components include Spatial-temporal Graph Convolution Networks (ST-GCN), Multi-layer perceptrons (MLP) and G-GRU networks. Details are in the supplementary.
@@ 3.1 A New Factorization of Interaction Motions
Directly learning $p(B_{s},q_{A},q_{B})$ would need large amounts of data containing different interactions with varying both lengths. Therefore, we first make it learnable on limited data by introducing a new factorization. First, we represent skeletons with different bone lengths as heterogeneously scaled versions of a template skeleton with a bone length scale vector $\hat{B}=\{1,...,1\}\in\mathbb{R}^{n}$, i.e. we treat the bone lengths of the template skeleton as scale 1. We abuse the notation and denote a skeleton variation by $B_{s}$, indicating how each bone is scaled with respect to $\hat{B}$.
@@ 3.2 Network Architecture
Instead of directly learning the distribution of $q^{\prime}_{B}$, learning the distribution of the differences $\triangle q_{B}=q^{\prime}_{B}-\hat{q}_{B}$ is easier [56, 49]: $p(q^{\prime}_{B}|B_{s},\hat{q}_{B})=p(\triangle q_{B}|B_{s})$, which is easier as it becomes learning the distribution of offsets from the template motion $\hat{q}_{B}$ and a skeleton variation $B_{s}$. We encode $\triangle q_{B}$ into a latent space then decode it back to the data space by:

### 09 Apr 2024 v2

@@ 3 Methodology
Figure 1: Overview of our model. The key components include Spatial-temporal Graph Convolution Networks (ST-GCN), Multi-layer perceptrons (MLP) and G-GRU networks. Details are in the supplementary material.
@@ 3.1 A New Factorization of Interaction Motions
Directly learning $p(B_{s},q_{A},q_{B})$ would need large amounts of data containing different interactions with varying bone lengths. Therefore, we first make it learnable on limited data by introducing a new factorization. First, we represent skeletons with different bone lengths as heterogeneously scaled versions of a template skeleton with a bone length scale vector $\hat{B}=\{1,...,1\}\in\mathbb{R}^{n}$, i.e. we treat the bone lengths of the template skeleton as scale 1. We abuse the notation and denote a skeleton variation by $B_{s}$, indicating how each bone is scaled with respect to $\hat{B}$.
@@ 3.2 Network Architecture
Instead of directly learning the distribution of $q^{\prime}_{B}$, learning the distribution of the differences $\triangle q_{B}=q^{\prime}_{B}-\hat{q}_{B}$ is easier [58, 51]: $p(q^{\prime}_{B}|B_{s},\hat{q}_{B})=p(\triangle q_{B}|B_{s})$, which is easier as it becomes learning the distribution of offsets from the template motion $\hat{q}_{B}$ and a skeleton variation $B_{s}$. We encode $\triangle q_{B}$ into a latent space and then decode it back to the data space by: